OpenStudy (anonymous):

Chose the expression that forms an equation with the given expression 9x2+10 divided by 2

OpenStudy (yanasidlinskiy):

Is that \(\huge{9*2??}\) or \(\huge{9x^2}\)

OpenStudy (anonymous):

It's 9 times 2

OpenStudy (yanasidlinskiy):

\(\huge\color{red}{9*2=18}\) \(\huge\color{blue}{18+10=28}\) \(\huge\color{black}{28/2 =14 }\) Is that what you need? Can you list your answer choices?

OpenStudy (anonymous):

So it's 20 divide 4 + 4 times 4